### v1

# Optimizing Allreduce Operations for Heterogeneous Architectures with Multiple Processes per GPU

Abstract.
Large inter-GPU all-reduce operations, prevalent throughout deep learning, are bottlenecked by communication costs. Emerging heterogeneous architectures are comprised of complex nodes, often containing $4$ GPUs and dozens to hundreds of CPU cores per node. Parallel applications are typically accelerated on the available GPUs, using only a single CPU core per GPU while the remaining cores sit idle. This paper presents novel optimizations to large GPU-aware all-reduce operations, extending lane-aware reductions to the GPUs, and notably using multiple CPU cores per GPU to accelerate these operations. These multi-CPU-accelerated GPU-aware lane all-reduces yield speedup of up to 2.452.45x for large MPI all-reduces across the NVIDIA A100 GPUs of NCSA’s Delta supercomputer. Finally, the approach is extended to NVIDIA’s and AMD’s collective communication libraries, achieving speedup of up to 1.771.77x and 1.711.71x, respectively, across $2$ state-of-the-art supercomputers.
Allreduce, heterogeneous, GPU, MPI, NCCL, RCCL, multi-lane, node-aware
$c$s: Computing methodologies Parallel algorithmsccs: Computing methodologies Massively parallel algorithmsccs: Computing methodologies Distributed algorithmsccs: Computer systems organization Parallel architectures
1. Introduction
Emerging exascale systems are comprised of heterogeneous nodes, each typically containing 484-8 GPUs and dozens to hundreds of CPU cores. Parallel applications achieve high performance through device utilization, accelerating computation on the available GPUs and communicating directly between devices with GPU-aware MPI. Applications typically use a single MPI process per GPU, utilizing one core optimally located near each GPU. While these architectures are equipped with dozens of CPU cores per node, the vast majority of these cores are unused by parallel applications.
Large all-reduce operations dominate the cost of many popular applications, including moderately sized neural networks and large language models (LLMs), in which the layers are distributed across GPUs.
Figure 1. Example All-Reduce Setup
[…]
Example buffers to be reduced across $8$ GPUs and $2$ nodes.
These reductions require each process to exchange and reduce data with all other processes. For example, Figure 1 shows input buffers across $8$ GPUs that are to be reduced.
While there are many existing all-reduce algorithms, the standard approach for large data sizes is through a ring pattern (ring0, ), in which every process exchanges a portion of the data with neighbors at each step, the first of which is displayed in Figure 2.
Figure 2. Ring All-Reduce, First Step
[…]
The first step of a ring all-reduce.
The cost of each message varies with the relative locations of the sending and receiving processes (perf_model_locality, ), and the ring algorithm attempts to optimize locality by restricting communication to neighboring ranks. Assuming an optimal process layout, the majority of processes may communicate locally, such as within a node. For example, the majority of messages displayed in Figure 2 are intra-node. However, at the boundary, processes must communicate with neighboring nodes at each step.
The multi-lane all-reduce (lane, ) improves the cost of large reductions through lane-awareness, with each process per node performing a separate portion of the inter-node all-reduce, exemplified in Figure 3.
Figure 3. Lane All-Reduce, Inter-Node Step
[…]
The inter-node step of a multi-lane all-reduce.
All processes first perform a reduce-scatter on-node, evenly partitioning the buffers before inter-node communication. Each process then exchanges smaller buffers with the corresponding processes on each other node. Finally, data is gathered within each node. While more data is communicated than with the ring algorithm, the additional overheads are within each node.
This paper explores methods for optimizing large GPU-aware all-reduce operations for emerging heterogeneous architectures. This paper presents a novel extension of the multi-lane algorithm to GPU-aware all-reduces. The authors further optimize the all-reduce through improved resource utilization, with multiple CPU cores each progressing a portion of the all-reduce asynchronously. Finally, the authors apply an equivalent technique to the NVIDIA and AMD collective communication libraries (CCLs) (NVIDIA2025NCCL, ; AMD2025RCCL, ), using multiple processes per GPU to accelerate the widely used CCL all-reduce.
Figure 4. Standard MPI_Allreduce vs our optimized all-reduce
[…]
This paper provides the following novel contributions:
  • Compares the performance of GPU-aware standard and multi-lane algorithms;
  • Uses multiple CPU cores per GPU, each progressing a portion of the all-reduce asynchronously;
  • Presents a performance analysis of these optimizations, achieving up to 2.452.45x speedup over OpenMPI’s GPU-aware all-reduce; and
  • Accelerates NCCL and RCCL all-reduce operations by up to 1.771.77x and 1.711.71x, respectively, through the use of multiple processes per GPU.
The remainder of our paper is outlined as follows. Section 2 describes background information and related works, detailing the baseline ring and multi-lane all-reduce algorithms. The methodology for optimizing large all-reduces is described in Section 3. Section 4 provides performance results for our approaches, with MPI results shown in Section 4.1 and extensions to the CCLs shown in Section 4.2. Finally, Section 5 provides concluding remarks.
2. Background
Input: rank, count, size from communicator
Get $r$, $c_{buf}$, $n$ from communicator
$c_{chunk}=c_{buf}/n$
$c_{onchunk}[n]$
for $i\leftarrow 0$ to $n$ do
 $c_{onchunk}[i]=c_{chunk}$
// displacements
$D[n]$ $D[0]=0$ for $i\leftarrow 1$ to $n$ do
 $D[i]=D[i-1]+c_{onchunk}[i-1]$
$sp=r$, rp=r1rp=r-1 if $r\neq 0$ else n1n-1
$buf_{sendfrom}=buf_{send}$
for $i\leftarrow 0$ to n1n-1 do
 // order based on rank
   MPI_Reduce to next process in ring of $c_{onchunk}[sp]$ counts of $buf_{sendfrom}$ at offset $D[sp]$
  MPI_Reduce from previous process in ring of $c_{onchunk}[rp]$ counts of $buf_{send}$ at offset $D[rp]$ to $buf_{recv}$ at offset $D[rp]$
 sp=sp1sp=sp-1 if $sp\neq 0$ else n1n-1
 rp=rp1rp=rp-1 if $rp\neq 0$ else n1n-1
 $buf_{sendfrom}=buf_{recv}$
$sp=r+1$ if $r\neq n-1$ else 0, $rp=r$
for $i\leftarrow 0$ to n1n-1 do
   MPI_Isend to next process in ring of $c_{onchunk}[sp]$ counts of $buf_{recv}$ at offset $D[sp]$
  MPI_Irecv from previous process in ring of $c_{onchunk}[rp]$ counts of $buf_{recv}$ at offset $D[rp]$
  MPI_Waitall
 sp=sp1sp=sp-1 if $sp\neq 0$ else n1n-1
 rp=rp1rp=rp-1 if $rp\neq 0$ else n1n-1
Algorithm 1 ring_allreduce
All-reduce operations are generally performed using three different algorithms: recursive doubling (recursive_doubling, ), Rabenseifner’s algorithm (ring0, ), and ring (ring0, ). Recursive doubling is used for small reductions, minimizing the number of messages to $\log{p}$ where $p$ is the number of processes, but sending all $n$ bytes of the buffer in each message. Rabenseifner’s algorithm improves performance for moderately-sized buffers, with each process reducing a fraction of the buffer before gathering all reductions. The total number of messages increases to $2\cdot\log{p}$, but the amount of data communicated and reduced upon is cut in half at each step. The ring all-reduce further optimizes for large reductions by restricting communication only to immediate neighbors, increasing the message count to $2\cdot(p-1)$ (ring0, ; ring1, ). In recent years, this algorithm has been adopted within deep learning applications (ring2, ; ring3, ).
The ring implementation for large all-reduce operations, detailed in Algorithm 1, requires each process to send a fraction of the buffer to one neighbor n1n-1 times, while receiving a separate fraction of the buffer from its other neighbor. The received message is then reduced with the local buffer, before the process is repeated with the next successive fractions of the buffer. Finally, in a similar manner, these fractions are gathered to all ranks by communicating with a process’s neighbors.
While the ring algorithm improves the locality of communication over other tree-based approaches, there remains a large variability among these neighboring messages. Assuming there are $4$ processes per node, two processes on each node will communicate only intra-node messages, while the other two will either send to or receive from a neighboring node.
Input: Rank, count, size from group communicator
Get $r$, $c_{buf}$, $n_{group}$ from communicator
$comm_{group}$
$comm_{lane}$
$c_{group}=c_{buf}/n_{group}$
$c_{ongroup}[n_{group}]$ $D[n_{group}]$ $D[0]=0$
for $i\leftarrow 0$ to $n_{group}$ do
 $c_{ongroup}[i]=c_{group}$
for $i\leftarrow 1$ to $n_{group}$ do
 $D[i]=D[i-1]+c_{ongroup}[i-1]$
MPI_Reduce_scatter from $buf_{send}$ to $buf_{recv}$ at receive offset $r(c_{group})$ of $c_{ongroup}$ counts on $comm_{group}$
MPI_Allreduce to $buf_{recv}$ at offset $r(c_{group})$ of $c_{ongroup}[r]$ counts on $comm_{lane}$
MPI_Allgatherv to $buf_{recv}$ of $c_{ongroup}$ counts on $comm_{group}$ using $D$
Algorithm 2 lane_allreduce
Multi-lane collectives (lane, ) reduce the variability between message costs through node-awareness. The multi-lane all-reduce, detailed in Algorithm 2, first performs a reduce-scatter on-node so that each process within the node holds a fraction of the buffer. Then, each process performs an all-reduce with its corresponding process on each other node. Finally, the reduced buffer portions are gathered among all processes within a node. Each process sending an equal portion of the buffer allows for the operation to better reach injection bandwidth limits since per-process inter-node messages are reduced in both size and count. Multi-lane algorithms have shown to greatly improve the performance of collectives on many-core architectures and heterogeneous systems requiring copies of the device buffer to host memory (gpu_c2c_lane, ). This paper presents a novel extension and analysis of these algorithms to GPU-aware all-reduce operations on heterogeneous systems using remote direct memory access (RDMA) communication.
2.1. GPU-Aware All-Reduce Operations
GPU-aware MPI implementations allow the CPU to offload communication to the GPU and network interface card (NIC) by utilizing the GPUDirect RDMA protocol. The NICs on emerging supercomputers support RDMA, allowing data to flow directly between the NIC and the GPU, bypassing copies with host memory. Large GPU-aware all-reduce operations take advantage of GPUDirect RDMA. While manually copying data to the CPU showed benefits on Power9 systems (perf_hetero, ), more modern systems obtain the best performance through the use of GPUDirect RDMA communication.
Parallel applications that rely on large all-reduces typically utilize a single MPI process per GPU. To achieve high-performance within the all-reduce, each process performs a GPU-aware all-reduce on the buffer, utilizing GPUDirect RDMA communication at each step of the algorithm. This requires the entire buffer to be communicated before a kernel is launched to perform the local reduction. The buffer is then sent on through the next step of the all-reduce only after the entire reduction is complete.
On many-core CPU systems, similar synchronization overheads have been reduced through partitioned communication, which allows each thread to instantiate communication on a local portion of the buffer without waiting for all threads to finish computation (partitioned, ). Similarly, partitioned communication allows for portions of the buffer to be received early, so that each thread can begin local computation as soon as its portion of the receive arrives, rather than waiting for the full buffer. While partitioned communication has been shown to optimize many operations on CPU-only systems, there have been extensions to heterogeneous systems. Due to the high overheads of kernel launches, along with the inability to initiate communication directly from the GPU, partitioning heterogeneous data into smaller partitions can incur bottlenecks.
While GPUDirect RDMA offloads some of the communication cost to the GPU, it is still highly dependent on CPU and NIC performance. Thus, this paper presents a novel optimization for MPI all-reduce operations to share a single GPU across multiple cores of the system, with each advancing a separate portion of the buffer. Ideally, this increases the bandwidth utilization of the NICs and Non-uniform memory access (NUMA) region interconnects while reducing synchronization overheads, allowing for each process to progress as soon as its portion of the buffer is ready.
2.2. The CCLs
Proprietary communication libraries, such as NVIDIA’s Collective Communications Library (NCCL) and AMD’s ROCm Communication Collectives Library (RCCL), implement the all-reduce with increased knowledge about the GPU and NIC hardware in use, yielding optimizations over MPI implementations that make them commonplace in AI applications. These all-reduce operations achieve very high bandwidth, optimal for large reductions. Similar to MPI, these all-reduces are typically performed with a single process per GPU. This paper presents a novel optimization to CCL all-reduces in which multiple processes per GPU each initiate a separate CCL all-reduce on a unique subset of the buffer.
2.3. Related Work
Other collectives such as the all-to-all, all-gather and related neighborhood methods feature a similar communication pattern to the all-reduce where in the worst case, every pair of processes must communicate. Intra-node communication has been optimized with shared memory (shared_memory, ; GrahamSharedMemoryColl2008, ; JainSharedMemoryColl2018, ) on CPUs. Hierarchical collectives aggregate all data onto a single leader per node, reducing inter-node traffic (hierarchical, ; KaronisHierColl2000, ; GrahamHierCheetah2011, ; TraffHierAllgather2006, ). Multi-leader collectives perform hierarchical algorithms with multiple leaders per node (multileader_hierarchical, ; multileader_hierarchical2, ). Locality-aware aggregation has also been used to aggregate messages, reducing the number of inter-node steps, both in the context of small all-reduce operations (bienz_allreduce, ), as well as other collective operations (bienz_bruck_allgather, ; locality_aware_alltoall, ).
Topology-aware approaches for collective operations have also been developed. For instance, for a cluster layout consisting of a two-dimensional grid of nodes where each node is directly connected to its four neighboring nodes, an all-reduce can be performed twice, once in each axis direction (BhateleTorusColl2012, ; torus, ), reducing the number of messages from the total number of nodes to the sum of all axes dimensions. Topology-aware mapping is commonly used to map processes within other topologies (topo0, ; topo1, ; topo2, ), such as trees (PatarasukTreeAllreduce2007, ), to minimize long-distance messages.
GPU-aware optimizations for collectives have been widely explored in recent years. Hierarchical methods have been analyzed on heterogeneous architectures (infini_hierarchical, ; gpu_multileader_hierarchical, ; hierarchical_ccl, ), communication primitives and collectives have been optimized to use direct memory access between GPUs more efficiently (gpu_dma, ; gpu_dma2, ; gpu_dma3, ; gpu_dma4, ), and compressed collectives have been used to reduce communication costs (comp_cosmo, ; comp_alltoall, ; comp_allreduce, ; zccl, ).
3. Methods
((a)) 8 nodes, varying buffer sizes
((b)) $2^{27}$ floats, varying node counts
Figure 5. Multi-Lane GPU-Aware Performance
Baseline GPU-aware all-reduce operations implement standard algorithms, such as the ring and multi-lane implementations in Algorithms 1 and 2, using GPUDirect RDMA for all communication. Further, these baseline implementations use a single MPI process per GPU.
3.1. Multiple Processes Per GPU
The baseline all-reduce operations are further optimized with multiple processes per GPU.
Figure 6. Partitioning Buffer Over Multiple Processes Per GPU
[…]
Each GPU on a node is assigned to an equal number of unique processes, as exemplified in Figure 6. In this example, there are two processes per GPU, displayed as red and blue outlines. The buffers to be reduced are partitioned evenly across the multiple processes per node.
Assuming there are PPG processes per GPU, each process $r$ on a node has a local rank $l_{r}\leftarrow r\mod{\texttt{PPG}}$. Each process with local rank $l_{r}=0$ is assigned as a leader process. The leader process creates the full buffer to be reduced on its corresponding device. After the buffer is created, the leader extracts its Inter-Process Communication (IPC) memory handle with (cuda/hip)GetIPCMemHandle. This memory handle is then broadcast to all processes assigned to the given GPU. Each non-leader process gains access to the shared device pointer by opening the received memory handle, using (cuda/hip)IpcOpenMemHandle.
Before an all-reduce operation can occur, new communicators must be created. A communicator, new_comm, must be created for all processes with equal local rank $l_{r}$, allowing for all-reduce operations on each communicator to reduce equivalent fractions of the buffer among all GPUs. All processes with local rank $l_{r}$ operate on the same fraction of a buffer located at $buf[s*l_{r}]$, but are assigned to separate GPUs. For example, all processes outlined as red in Figure 6 will reduce the first half of all buffers, while the blue processes reduce the second half.
3.1.1. Standard Approach
After the initial setup, each process performs an equal portion of the all-reduce asynchronously. Processes assigned to the same offset in the buffer of each GPU perform an all-reduce. For example, two all-reduces are performed concurrently in Figure 6; one among all red processes, and a second among all blue processes.
In the baseline case with only a single process per GPU, all processes would reduce a buffer of $s$ bytes with the following.
MPI_Allreduce(sendbuf, recvbuf, s, MPI_SUM,
MPI_COMM_WORLD);
When using multiple processes per GPU, this all-reduce call becomes the following.
MPI_Allreduce(&(sendbuf[s*l_r]),
&(recvbuf[s*l_r]),
s/PPG, MPI_SUM, new_comm);
In this algorithm, the total number of messages among all processes increases by the factor of the number of processes per GPU. However, each message is reduced in size by the same factor. These messages are communicated concurrently.
Figure 7. Multi-PPG, Multi-Lane
[…]
The number of kernel launches per process remains constant. As the MPI all-reduce operation does not allow passing streams and the NVIDIA Multi-Process Service (MPS) is assumed to be disabled, the local reductions are unable to be performed in parallel. However, as there is asynchrony among processes, all kernels are unlikely to compete for the GPU at one time.
3.1.2. Multi-lane Approach
The standard approach for using multiple processes per GPU can be further extended to utilize multi-lane collectives. The multi-PPG all-reduce is broken into three stages: 1) intra-node reduce-scatter, 2) inter-node all-reduce on a lane communicator, and 3) intra-node all-gatherv, as described in Algorithm 2. In the multi-process version of the multi-lane algorithm, the send buffer, $buf_{send}$, and receive buffer, $buf_{recv}$, are passed in at offset $s*l_{r}$. Further, $comm_{group}$ contains corresponding processes of on node GPUs and $comm_{lane}$ contains corresponding processes on other nodes. To create these communicators, new_comm is split into processes that reside on the same node ($comm_{group}$) and processes that reside on separate nodes ($comm_{lane}$). With multiple processes per GPU, each CPU core is a separate lane.
The inter-node stage of this all-reduce is displayed in Figure 7. Initially, all red processes perform intra-node reduce-scatters. At the same time, blue processes perform separate intra-node reduce-scatters. After the scattered buffers are reduced, each process performs inter-node all-reduce operations with corresponding processes. Finally, red and blue processes concurrently gather their reduced data on-node.
The reduce-scatter divides the input buffer into chunks equal to the number of GPUs on a node. The off-node all-reduce messages are reduced in size proportionally, leading to a total reduction in inter-node buffer size by the number of MPI processes per node. However, there is an increase in intra-node communication from the reduce-scatter and the subsequent all-gatherv operations.
4. Results
((a)) 8 nodes, varying buffer sizes
((b)) $2^{27}$ floats, varying node counts
Figure 8. Standard and multi-lane (with MPI_Allreduce and ring all-reduce) algorithms with the maximum of multiple-processes per GPU
The impact of multi-lane and multi-process per GPU all-reduce algorithms is analyzed against OpenMPI, NCCL, and RCCL on state-of-the-art supercomputers: NCSA’s Delta and DeltaAI as well as LLNL’s Tuolumne. Some details of these systems are provided below.
Delta: This system provides a variety of nodes, including both NVIDIA and AMD GPUs. However, due to the limited number of AMD nodes, only NVIDIA nodes were considered for the scaling studies in this paper. All-reduce operations were analyzed on the single-socket nodes, each composed of a 64 core AMD Milan CPU with 4 NVIDIA A100 GPUs. DeltaAI: This supercomputer is comprised of Grace Hopper superchips. Each node is equipped with four NVIDIA GH200 superchips, each containing a Grace ARM 72-core CPU and a NVIDIA H100 GPU. Tuolumne: Each node has four AMD MI300A APUs, each of which has a 24-core AMD EPYC CPU and a CDNA 3 GPU.
The impact on MPI all-reduce operations is analyzed solely on Delta, as OpenMPI was available only on this system. All other systems provide Cray MPICH, which forbids any buffer from being communicated through GPU-aware MPI operations if it was not allocated by the calling process. As a result, sharing IPC buffers within GPU-aware MPI is not possible on these systems. The impact of multiple processes per GPU on the CCL libraries, however, was evaluated on Cray systems.
Note that NVIDIA MPS was not enabled during benchmarking.
4.1. MPI Analysis
All MPI all-reduce tests were analyzed using OpenMPI 5.0.5 and CUDA 12.4.0. All processes were assigned to the GPU adjacent to their NUMA region. Each benchmark was run $10$ separate times to demonstrate reproducibility. All plots contain error bars showing the run to run variation of the algorithms.
Figure 5 compares the performance of a standard MPI_Allreduce operation against the multi-lane optimization. These tests are using a single process per GPU, and using all four GPUs available per node on Delta. While the standard algorithm outperforms multi-lane at the smallest data sizes, the multi-lane algorithm yields significant speedups for larger all-reduces. The speedups for the multi-lane algorithm increase with node count, achieving 2.12.1x speedups at $32$ nodes.
Figure 9. A standard MPI_Allreduce with multiple processes per GPU.
Figure 9 shows the performance of a standard MPI_Allreduce when varying the number of processes per GPU from $1$ to $16$. Performance continually improves as the number of processes per GPU increases. Further, these speedups increase with node count through $16$ nodes, yielding over 1.71.7x speedup at $32$ nodes.
Figure 10. The multi-lane all-reduce with multiple processes per GPU.
((a)) Tuolumne (RCCL)
((b)) DeltaAI (NCCL)
Figure 11. Standard CCL algorithm with a size of $2^{29}$ floats, varying node counts
Figure 10 displays the performance of the multi-lane algorithm when using multiple processes per GPU. Similar to the standard MPI_Allreduce, the multi-lane algorithm is further accelerated with increasing numbers of processes per GPU. Large all-reduce operations across all $4$ GPUs on $8$ nodes yield around a 1.391.39x speedup over the multi-lane algorithm with a single process per GPU.
Finally, Figure 8 compares the performance of all algorithms, including where the ring algorithm is used in the inter-node stage of the multi-lane all-reduce, comparing $1$ versus $16$ processes per GPU. The multi-lane algorithm with $16$ processes per GPU achieves up to 2.42.4x speedup over the standard MPI_Allreduce for large data sizes.
4.2. CCL Analysis
This section analyzes the impact of using multiple processes per GPU across CCL all-reduce operations. The performance of NCCL is analyzed on a state-of-the-art NVIDIA Grace Hopper architecture, Delta AI, using CUDA 12.3 and NCCL 2.19.3. The performance of RCCL is analyzed on one of the fastest AMD architectures, Tuolumne, using ROCm 6.4.0 and RCCL 2.22.3. Each benchmark was run 3 separate times; error bars in plots show run to run variation.
Each process performs an independent CCL all-reduce operation on their assigned subset of the buffer. As CCL all-reduce operations take a stream, each process per GPU performs this reduction on a separate stream to allow increased parallelism. All processes were assigned to the GPU adjacent to their NUMA region.
Figure 11 shows the performance of a CCL all-reduce at the largest size when varying the number of processes per GPU from 1 to 16 on Tuolumne and DeltaAI. On the left featuring Tuolume using RCCL, we see steady speedup as the number of nodes increases with 2 and 4 processes per GPU, yielding speedups of 1.711.71x and 1.591.59x at 8 nodes, respectively. These speedups begin to decrease at large scales. Using more than 4 processes per GPU does not show any benefit on Tuolumne past 8 nodes with 8 and 16 processes per GPU showing significant slowdown.
On the right featuring DeltaAI using NCCL, we see substantial speedup at all node counts with 16 processes per GPU, yielding the most significant speedup of 1.771.77x at 16 nodes. 4 and 8 processes per GPU show speedup only at small scales, while using 2 processes per GPU starts to show benefit at 16 nodes.
There are many potential reasons for this speedup, including improved network utilization and improved use of NCCL channels. To determine if NCCL channels were impacted, the environment variables NCCL_MIN_NCHANNELS and NCCL_MAX_NCHANNELS were set to 4 in order to eliminate the case of low GPU core utilization during memory operations; no differences were observed in timings.
5. Conclusions and Future Work
In this paper, we proposed to utilize more of the available CPU cores across all GPUs on a node to speed up the GPU-aware all-reduce collective. Using an Ampere-based system (Delta) on OpenMPI with our standard and lane-aware approaches, we showed up to an 2.452.45x speedup on large buffers. Using RCCL with our standard approach on an El Capitan-based system, we showed a 1.711.71x speedup. Using NCCL on a Hopper-based system, we showed a 1.771.77x speedup. We conclude that using more of the available CPU cores on modern systems should enhance all-reduce performance at large scales.
We propose to extensively evaluate the performance characteristics of our multi-process approaches and also extend them to collectives such as the neighborhood all-to-all used in sparse linear systems. Due to the hierarchical structure of modern clusters, future work may also extend this approach to cluster topologies, i.e. stages of reduction in fat tree networks, in order to reduce the number of processes required at each stage of the communication.
Emerging systems like Tuolumne feature memory architectures that are shared between the CPU and GPU. Future work could potentially examine tradeoffs using CPUs for communication and GPUs for computation, allowing these steps to overlap while minimizing synchronization costs.
Acknowledgments
This work was performed with partial support from the National Science Foundation under Grant No. CCF-2338077, the U.S. Department of Energy’s National Nuclear Security Administration (NNSA) under the Predictive Science Academic Alliance Program (PSAAP-III), Award DE-NA0003966.
Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation and the U.S. Department of Energy’s National Nuclear Security Administration.
This research used the Tuolumne supercomputer at Lawrence Livermore National Laboratory. This research used both the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
Generative AI was used to improve the formatting of data within plots.

### v2

@@
-Optimizing Allreduce Operations for Heterogeneous Architectures with Multiple Processes per GPU
-Abstract.
+Optimizing Allreduce Operations for Modern Heterogeneous Architectures with Multiple Processes per GPU
+Abstract
 Large inter-GPU all-reduce operations, prevalent throughout deep learning, are bottlenecked by communication costs.
@@
 Parallel applications are typically accelerated on the available GPUs, using only a single CPU core per GPU while the remaining cores sit idle.
-This paper presents novel optimizations to large GPU-aware all-reduce operations, extending lane-aware reductions to the GPUs, and notably using multiple CPU cores per GPU to accelerate these operations.
-These multi-CPU-accelerated GPU-aware lane all-reduces yield speedup of up to 2.452.45x for large MPI all-reduces across the NVIDIA A100 GPUs of NCSA’s Delta supercomputer.
-Finally, the approach is extended to NVIDIA’s and AMD’s collective communication libraries, achieving speedup of up to 1.771.77x and 1.711.71x, respectively, across $2$ state-of-the-art supercomputers.
-Allreduce, heterogeneous, GPU, MPI, NCCL, RCCL, multi-lane, node-aware
-$c$s: Computing methodologies Parallel algorithmsccs: Computing methodologies Massively parallel algorithmsccs: Computing methodologies Distributed algorithmsccs: Computer systems organization Parallel architectures
-1. Introduction
+This paper presents novel optimizations to large GPU-aware all-reduce operations by extending the lane-aware algorithm to heterogeneous architectures and notably using multiple CPU cores per GPU to accelerate these operations.
+Using GPUDirect RDMA and host copy communications respectively, these multi-CPU-accelerated GPU-aware all-reduces yield speedups over system MPI of up to 33x on LLNL’s Tuolumne supercomputer and up to 2.452.45x for large MPI all-reduces across the NVIDIA A100 GPUs of NCSA’s Delta supercomputer.
+I Introduction
 Emerging exascale systems are comprised of heterogeneous nodes, each typically containing 484-8 GPUs and dozens to hundreds of CPU cores.
@@
 Large all-reduce operations dominate the cost of many popular applications, including moderately sized neural networks and large language models (LLMs), in which the layers are distributed across GPUs.
-Figure 1. Example All-Reduce Setup
-[…]
-Example buffers to be reduced across $8$ GPUs and $2$ nodes.
+Figure 1: Example All-Reduce Setup
 These reductions require each process to exchange and reduce data with all other processes. For example, Figure 1 shows input buffers across $8$ GPUs that are to be reduced.
-While there are many existing all-reduce algorithms, the standard approach for large data sizes is through a ring pattern (ring0, ), in which every process exchanges a portion of the data with neighbors at each step, the first of which is displayed in Figure 2.
-Figure 2. Ring All-Reduce, First Step
-[…]
-The first step of a ring all-reduce.
-The cost of each message varies with the relative locations of the sending and receiving processes (perf_model_locality, ), and the ring algorithm attempts to optimize locality by restricting communication to neighboring ranks.
+While there are many existing all-reduce algorithms, the standard approach for large data sizes is through a ring pattern [35], in which every process exchanges a portion of the data with neighbors at each step, the first of which is displayed in Figure 2.
+Figure 2: Ring All-Reduce, First Step
+The cost of each message varies with the relative locations of the sending and receiving processes [7], and the ring algorithm attempts to optimize locality by restricting communication to neighboring ranks.
 Assuming an optimal process layout, the majority of processes may communicate locally, such as within a node.
@@
 However, at the boundary, processes must communicate with neighboring nodes at each step.
-The multi-lane all-reduce (lane, ) improves the cost of large reductions through lane-awareness, with each process per node performing a separate portion of the inter-node all-reduce, exemplified in Figure 3.
-Figure 3. Lane All-Reduce, Inter-Node Step
-[…]
-The inter-node step of a multi-lane all-reduce.
+The multi-lane all-reduce [40] improves the cost of large reductions through lane-awareness, with each process per node performing a separate portion of the inter-node all-reduce, exemplified in Figure 3.
+Figure 3: Lane All-Reduce, Inter-Node Step
 All processes first perform a reduce-scatter on-node, evenly partitioning the buffers before inter-node communication.
@@
 This paper explores methods for optimizing large GPU-aware all-reduce operations for emerging heterogeneous architectures.
-This paper presents a novel extension of the multi-lane algorithm to GPU-aware all-reduces.
-The authors further optimize the all-reduce through improved resource utilization, with multiple CPU cores each progressing a portion of the all-reduce asynchronously.
-Finally, the authors apply an equivalent technique to the NVIDIA and AMD collective communication libraries (CCLs) (NVIDIA2025NCCL, ; AMD2025RCCL, ), using multiple processes per GPU to accelerate the widely used CCL all-reduce.
-Figure 4. Standard MPI_Allreduce vs our optimized all-reduce
-[…]
+This paper presents a novel extension of the multi-lane algorithm to GPU-aware all-reduces backed by host copy and GPUDirect Remote Direct Memory Access (RDMA) communication on NVIDIA Ampere-based and AMD MI300A-based systems, respectively.
+We further optimize the all-reduce through improved resource utilization, with multiple CPU cores each progressing a portion of the all-reduce asynchronously with performance as seen in Figure 4.
+Finally, we analyze the performance of the all-reduces in the various partitioning modes of the AMD MI300A, showing that the multi-lane all-reduces benefit from increased logical GPU count.
+Figure 4: Standard MPI_Allreduce vs our optimized all-reduce on LLNL’s Tuolumne
 This paper provides the following novel contributions:
-  • Compares the performance of GPU-aware standard and multi-lane algorithms;
+  • Compares the performance of GPU-aware standard and multi-lane algorithms on modern architectures;
   • Uses multiple CPU cores per GPU, each progressing a portion of the all-reduce asynchronously;
-  • Presents a performance analysis of these optimizations, achieving up to 2.452.45x speedup over OpenMPI’s GPU-aware all-reduce; and
-  • Accelerates NCCL and RCCL all-reduce operations by up to 1.771.77x and 1.711.71x, respectively, through the use of multiple processes per GPU.
+  • Presents a performance analysis of these optimizations on modern architectures, achieving up to 2.452.45x speedup over OpenMPI’s GPU-aware all-reduce using host copy communication and up to 1.171.17x speedup over Cray MPICH’s GPUDirect RDMA implementation; and
+  • Accelerates all-reduces using the TPX and CPX partitioning modes using the multi-lane algorithm, achieving up to 33x speedup over the Cray MPICH implementation.
 The remainder of our paper is outlined as follows.
-Section 2 describes background information and related works, detailing the baseline ring and multi-lane all-reduce algorithms.
-The methodology for optimizing large all-reduces is described in Section 3.
-Section 4 provides performance results for our approaches, with MPI results shown in Section 4.1 and extensions to the CCLs shown in Section 4.2.
-Finally, Section 5 provides concluding remarks.
-2. Background
+Section II describes background information and related works, detailing the baseline ring and multi-lane all-reduce algorithms.
+Section III provides benchmarks of point-to-point communication using multiple processes per GPU to motivate our work.
+The methodology for optimizing large all-reduces is described in Section IV.
+Section V provides performance results for our approaches, with results using host copy communication shown in Section V-A and extensions to GPUDirect RDMA communication on the AMD MI300A shown in Section V-B.
+Performance on the various partitioning modes of the AMD MI300A is shown in Section V-B1.
+Finally, Section VI provides concluding remarks.
+II Background
+All-reduce operations are generally performed using three different algorithms: recursive doubling [38], Rabenseifner’s algorithm [35], and ring [35].
+Recursive doubling is used for small reductions, minimizing the number of messages to $\log{p}$ where $p$ is the number of processes, but sending all $n$ bytes of the buffer in each message.
+Rabenseifner’s algorithm improves performance for moderately-sized buffers, with each process reducing a fraction of the buffer before gathering all reductions.
+The total number of messages increases to $2\cdot\log{p}$, but the amount of data communicated and reduced upon is cut in half at each step.
+The ring all-reduce further optimizes for large reductions by restricting communication only to immediate neighbors, increasing the message count to $2\cdot(p-1)$ [35, 33].
+In recent years, this algorithm has been adopted within deep learning applications [15, 37].
 Input: rank, count, size from communicator
 Get $r$, $c_{buf}$, $n$ from communicator
+1ex
 $c_{chunk}=c_{buf}/n$
@@
  $c_{onchunk}[i]=c_{chunk}$
+1ex
 // displacements
@@
 Algorithm 1 ring_allreduce
-All-reduce operations are generally performed using three different algorithms: recursive doubling (recursive_doubling, ), Rabenseifner’s algorithm (ring0, ), and ring (ring0, ).
-Recursive doubling is used for small reductions, minimizing the number of messages to $\log{p}$ where $p$ is the number of processes, but sending all $n$ bytes of the buffer in each message.
-Rabenseifner’s algorithm improves performance for moderately-sized buffers, with each process reducing a fraction of the buffer before gathering all reductions.
-The total number of messages increases to $2\cdot\log{p}$, but the amount of data communicated and reduced upon is cut in half at each step.
-The ring all-reduce further optimizes for large reductions by restricting communication only to immediate neighbors, increasing the message count to $2\cdot(p-1)$ (ring0, ; ring1, ).
-In recent years, this algorithm has been adopted within deep learning applications (ring2, ; ring3, ).
 The ring implementation for large all-reduce operations, detailed in Algorithm 1, requires each process to send a fraction of the buffer to one neighbor n1n-1 times, while receiving a separate fraction of the buffer from its other neighbor.
@@
 Assuming there are $4$ processes per node, two processes on each node will communicate only intra-node messages, while the other two will either send to or receive from a neighboring node.
-Input: Rank, count, size from group communicator
+Input: rank, count,
+   size from group communicator
 Get $r$, $c_{buf}$, $n_{group}$ from communicator
@@
 MPI_Reduce_scatter from $buf_{send}$ to $buf_{recv}$ at receive offset $r(c_{group})$ of $c_{ongroup}$ counts on $comm_{group}$
+1ex
 MPI_Allreduce to $buf_{recv}$ at offset $r(c_{group})$ of $c_{ongroup}[r]$ counts on $comm_{lane}$
@@
 Algorithm 2 lane_allreduce
-Multi-lane collectives (lane, ) reduce the variability between message costs through node-awareness.
+Multi-lane collectives [40] reduce the variability between message costs through node-awareness.
 The multi-lane all-reduce, detailed in Algorithm 2, first performs a reduce-scatter on-node so that each process within the node holds a fraction of the buffer.
@@
 Each process sending an equal portion of the buffer allows for the operation to better reach injection bandwidth limits since per-process inter-node messages are reduced in both size and count.
-Multi-lane algorithms have shown to greatly improve the performance of collectives on many-core architectures and heterogeneous systems requiring copies of the device buffer to host memory (gpu_c2c_lane, ).
-This paper presents a novel extension and analysis of these algorithms to GPU-aware all-reduce operations on heterogeneous systems using remote direct memory access (RDMA) communication.
-2.1. GPU-Aware All-Reduce Operations
+Multi-lane algorithms have shown to greatly improve the performance of collectives on older many-core architectures and heterogeneous systems requiring copies of the device buffer to host memory [13].
+This paper presents a novel extension and analysis of these algorithms to GPU-aware all-reduce operations on heterogeneous systems using GPUDirect RDMA communication; this paper also justifies their continued applicability on modern systems requiring copies between host and device memory.
+II-A GPU-Aware All-Reduce Operations
 GPU-aware MPI implementations allow the CPU to offload communication to the GPU and network interface card (NIC) by utilizing the GPUDirect RDMA protocol.
@@
 Large GPU-aware all-reduce operations take advantage of GPUDirect RDMA.
-While manually copying data to the CPU showed benefits on Power9 systems (perf_hetero, ), more modern systems obtain the best performance through the use of GPUDirect RDMA communication.
+While manually copying data to the CPU showed benefits on Power9 systems [9], more modern systems obtain the best performance through the use of GPUDirect RDMA communication.
 Parallel applications that rely on large all-reduces typically utilize a single MPI process per GPU.
@@
 The buffer is then sent on through the next step of the all-reduce only after the entire reduction is complete.
-On many-core CPU systems, similar synchronization overheads have been reduced through partitioned communication, which allows each thread to instantiate communication on a local portion of the buffer without waiting for all threads to finish computation (partitioned, ).
+On many-core CPU systems, similar synchronization overheads have been reduced through partitioned communication, which allows each thread to start communication on a local portion of the buffer without waiting for all threads to finish computation [19].
 Similarly, partitioned communication allows for portions of the buffer to be received early, so that each thread can begin local computation as soon as its portion of the receive arrives, rather than waiting for the full buffer.
-While partitioned communication has been shown to optimize many operations on CPU-only systems, there have been extensions to heterogeneous systems.
+While partitioned communication has been shown to optimize many operations on CPU-only systems, there have been extensions to heterogeneous systems [13].
 Due to the high overheads of kernel launches, along with the inability to initiate communication directly from the GPU, partitioning heterogeneous data into smaller partitions can incur bottlenecks.
+(a) (Delta)
+(b) Default SPX mode (Tuolumne)
+(c) Fine-grained CPX mode (Tuolumne)
+Figure 5: A ping-pong benchmark using multiple active processes per (logical) GPU featuring performance on Delta (left), and the SPX (middle) and CPX (right) modes of the AMD MI300A on Tuolumne
 While GPUDirect RDMA offloads some of the communication cost to the GPU, it is still highly dependent on CPU and NIC performance.
@@
 Ideally, this increases the bandwidth utilization of the NICs and Non-uniform memory access (NUMA) region interconnects while reducing synchronization overheads, allowing for each process to progress as soon as its portion of the buffer is ready.
-2.2. The CCLs
-Proprietary communication libraries, such as NVIDIA’s Collective Communications Library (NCCL) and AMD’s ROCm Communication Collectives Library (RCCL), implement the all-reduce with increased knowledge about the GPU and NIC hardware in use, yielding optimizations over MPI implementations that make them commonplace in AI applications.
+II-B Partitioning Modes of the AMD MI300A
+The AMD MI300A Accelerated Processing Unit (APU) is a processor that places the CPU and GPU in a single package [1].
+There are three CPU chiplets each with 8-core AMD Zen4 CPUs and six Accelerator Complex Dies (XCDs) with 38 CDNA3-based GPU compute units each.
+128 GB of HBM3 memory is accessible to both CPU cores and GPU compute units backed by 256 MB of shared last level cache.
+At runtime, the XCDs can be grouped logically depending on the user’s needs.
+By default, the MI300A APU operates in Single Partition X-celerator (SPX) mode in which all 6 XCDs are exposed to the user as a single GPU.
+In this case, the workload on the GPU is automatically distributed across all XCDs.
+With the Triple Partition X-celerator (TPX) mode, the APU is treated as three GPUs of two XCDs each.
+In Core Partitioned X-celerator (CPX) mode, each XCD is exposed as a single GPU.
+The TPX and CPX partitioning modes offer the developer more granular control over the hardware that is chosen to execute their applications.
+Furthermore, since all partitioning modes can access the entire pool of HBM3 memory, these modes offer increased parallelism when used effectively.
+II-C Related Work
+Other collectives such as the all-to-all, all-gather and related neighborhood methods feature a similar communication pattern to the all-reduce where in the worst case, every pair of processes must communicate.
+Intra-node communication has been optimized with shared memory [20, 17, 23] on CPUs.
+Hierarchical collectives aggregate all data onto a single leader per node, reducing inter-node traffic [45, 27, 18, 41].
+Multi-leader collectives perform hierarchical algorithms with multiple leaders per node [26, 4].
+Locality-aware aggregation has also been used to aggregate messages, reducing the number of inter-node steps, both in the context of small all-reduce operations [8], as well as other collective operations [6, 10].
+Topology-aware approaches for collective operations have also been developed.
+For instance, for a cluster layout consisting of a two-dimensional grid of nodes where each node is directly connected to its four neighboring nodes, an all-reduce can be performed twice, once in each axis direction [5, 29], reducing the number of messages from the total number of nodes to the sum of all axes dimensions.
+Topology-aware mapping is commonly used to map processes within other topologies [12, 30, 28], such as trees [32], to minimize long-distance messages.
+GPU-aware optimizations for collectives have been widely explored in recent years.
+Hierarchical methods have been analyzed on heterogeneous architectures [16, 39, 21], communication primitives and collectives have been optimized to use direct memory access between GPUs more efficiently [24, 34, 42, 11], and compressed collectives have been used to reduce communication costs [25, 43, 44, 22].
+Proprietary communication libraries, such as NVIDIA’s Collective Communications Library (NCCL) [31] and AMD’s ROCm Communication Collectives Library (RCCL) [2], implement the all-reduce with increased knowledge about the GPU and NIC hardware in use, yielding optimizations over MPI implementations that make them commonplace in AI applications.
 These all-reduce operations achieve very high bandwidth, optimal for large reductions.
 Similar to MPI, these all-reduces are typically performed with a single process per GPU.
-This paper presents a novel optimization to CCL all-reduces in which multiple processes per GPU each initiate a separate CCL all-reduce on a unique subset of the buffer.
-2.3. Related Work
-Other collectives such as the all-to-all, all-gather and related neighborhood methods feature a similar communication pattern to the all-reduce where in the worst case, every pair of processes must communicate.
-Intra-node communication has been optimized with shared memory (shared_memory, ; GrahamSharedMemoryColl2008, ; JainSharedMemoryColl2018, ) on CPUs.
-Hierarchical collectives aggregate all data onto a single leader per node, reducing inter-node traffic (hierarchical, ; KaronisHierColl2000, ; GrahamHierCheetah2011, ; TraffHierAllgather2006, ).
-Multi-leader collectives perform hierarchical algorithms with multiple leaders per node (multileader_hierarchical, ; multileader_hierarchical2, ).
-Locality-aware aggregation has also been used to aggregate messages, reducing the number of inter-node steps, both in the context of small all-reduce operations (bienz_allreduce, ), as well as other collective operations (bienz_bruck_allgather, ; locality_aware_alltoall, ).
-Topology-aware approaches for collective operations have also been developed.
-For instance, for a cluster layout consisting of a two-dimensional grid of nodes where each node is directly connected to its four neighboring nodes, an all-reduce can be performed twice, once in each axis direction (BhateleTorusColl2012, ; torus, ), reducing the number of messages from the total number of nodes to the sum of all axes dimensions.
-Topology-aware mapping is commonly used to map processes within other topologies (topo0, ; topo1, ; topo2, ), such as trees (PatarasukTreeAllreduce2007, ), to minimize long-distance messages.
-GPU-aware optimizations for collectives have been widely explored in recent years.
-Hierarchical methods have been analyzed on heterogeneous architectures (infini_hierarchical, ; gpu_multileader_hierarchical, ; hierarchical_ccl, ), communication primitives and collectives have been optimized to use direct memory access between GPUs more efficiently (gpu_dma, ; gpu_dma2, ; gpu_dma3, ; gpu_dma4, ), and compressed collectives have been used to reduce communication costs (comp_cosmo, ; comp_alltoall, ; comp_allreduce, ; zccl, ).
-3. Methods
-((a)) 8 nodes, varying buffer sizes
-((b)) $2^{27}$ floats, varying node counts
-Figure 5. Multi-Lane GPU-Aware Performance
-Baseline GPU-aware all-reduce operations implement standard algorithms, such as the ring and multi-lane implementations in Algorithms 1 and 2, using GPUDirect RDMA for all communication.
+III Microbenchmarks
+GPU-aware operations, such as the all-reduce, are performed using one controlling process per GPU.
+This process initiates all GPU-aware communication involving its GPU.
+However, modern systems feature many more CPU cores than GPUs, causing them to be underutilized.
+To demonstrate the potential advantage of using multiple processes per GPU, we evaluated a ping-pong benchmark on both systems analyzed.
+In this benchmark, there are two nodes where each rank on a node sends and receives a message with its corresponding rank on the other node.
+Each GPU on a node is participating in the benchmark and controlled by at least 1 MPI process.
+To obtain the time it takes to send or receive one message, the round-trip time is divided by $2$.
+Benchmark results on the Delta system with intermediate copies between host and device memory can be seen in Figure 5(a).
+When 16 processes per GPU are used, this yields a significant speedup of 3.493.49x over a standard ping-pong of the largest message size, suggesting a benefit by the use of multiple processes per GPU during communication on Delta.
+Results using GPUDirect RDMA communication on the SPX and CPX partitioning modes of the APU on Tuolumne can be seen in Figures 5(b) and 5(c).
+Similarly to results on Delta, the ping-pong using SPX mode yields speedups of up to 1.891.89x over a standard ping-pong when using 16 processes per GPU.
+However, when fine-graned control of each XCD is enabled with CPX mode, using multiple processes per GPU does not show significant benefit.
+Therefore, these results suggest using mutiple processes per GPU is more applicable when using the coarse-grained partitioning on Tuolumne’s APUs.
+IV Methods
+Baseline GPU-aware all-reduce operations implement standard algorithms, such as the ring and multi-lane implementations in Algorithms 1 and 2.
 Further, these baseline implementations use a single MPI process per GPU.
-3.1. Multiple Processes Per GPU
+IV-A Multiple Processes Per GPU
 The baseline all-reduce operations are further optimized with multiple processes per GPU.
-Figure 6. Partitioning Buffer Over Multiple Processes Per GPU
-[…]
 Each GPU on a node is assigned to an equal number of unique processes, as exemplified in Figure 6.
@@
 For example, all processes outlined as red in Figure 6 will reduce the first half of all buffers, while the blue processes reduce the second half.
-3.1.1. Standard Approach
+IV-A1 Standard Approach
+Figure 6: Partitioning Buffer Over Multiple Processes Per GPU
 After the initial setup, each process performs an equal portion of the all-reduce asynchronously.
@@
 These messages are communicated concurrently.
-Figure 7. Multi-PPG, Multi-Lane
-[…]
 The number of kernel launches per process remains constant.
-As the MPI all-reduce operation does not allow passing streams and the NVIDIA Multi-Process Service (MPS) is assumed to be disabled, the local reductions are unable to be performed in parallel.
-However, as there is asynchrony among processes, all kernels are unlikely to compete for the GPU at one time.
-3.1.2. Multi-lane Approach
+The host copy implementations are likely to be performed in parallel excluding the host-device copy operations since the Multi-Process Service (MPS) is disabled on the Ampere-based system.
+For the GPUDirect RDMA implementations, communication is either parallelized using MPS or, as there is asynchrony among processes, all kernels are unlikely to compete for the GPU at one time.
+It is noted on the AMD MI300A system tested, AMD has natively supported MPS features in their GPU driver [14], so MPS is implied to be enabled.
+IV-A2 Multi-lane Approach
 The standard approach for using multiple processes per GPU can be further extended to utilize multi-lane collectives.
@@
 With multiple processes per GPU, each CPU core is a separate lane.
+Figure 7: Multi-PPG, Multi-Lane
 The inter-node stage of this all-reduce is displayed in Figure 7.
@@
 However, there is an increase in intra-node communication from the reduce-scatter and the subsequent all-gatherv operations.
-4. Results
-((a)) 8 nodes, varying buffer sizes
-((b)) $2^{27}$ floats, varying node counts
-Figure 8. Standard and multi-lane (with MPI_Allreduce and ring all-reduce) algorithms with the maximum of multiple-processes per GPU
-The impact of multi-lane and multi-process per GPU all-reduce algorithms is analyzed against OpenMPI, NCCL, and RCCL on state-of-the-art supercomputers: NCSA’s Delta and DeltaAI as well as LLNL’s Tuolumne.
+V Results
+The impact of multi-lane and multi-process per GPU all-reduce algorithms is analyzed against OpenMPI and Cray MPICH on state-of-the-art supercomputers: NCSA’s Delta and LLNL’s Tuolumne.
 Some details of these systems are provided below.
 Delta: This system provides a variety of nodes, including both NVIDIA and AMD GPUs.
-However, due to the limited number of AMD nodes, only NVIDIA nodes were considered for the scaling studies in this paper.
+The number of AMD nodes is limited, so only NVIDIA nodes were benchmarked for this paper.
 All-reduce operations were analyzed on the single-socket nodes, each composed of a 64 core AMD Milan CPU with 4 NVIDIA A100 GPUs.
-DeltaAI: This supercomputer is comprised of Grace Hopper superchips.
-Each node is equipped with four NVIDIA GH200 superchips, each containing a Grace ARM 72-core CPU and a NVIDIA H100 GPU.
 Tuolumne: Each node has four AMD MI300A APUs, each of which has a 24-core AMD EPYC CPU and a CDNA 3 GPU.
-The impact on MPI all-reduce operations is analyzed solely on Delta, as OpenMPI was available only on this system.
-All other systems provide Cray MPICH, which forbids any buffer from being communicated through GPU-aware MPI operations if it was not allocated by the calling process.
+The impact on MPI all-reduce operations using copies between host and device memory is analyzed solely on Delta, as only OpenMPI was available on this system at the time of writing.
+Tuolumne supports GPUDirect RDMA provided by Cray MPICH.
+Cray MPICH does not support any buffer from being communicated through GPU-aware MPI operations if it was not allocated by the calling process.
 As a result, sharing IPC buffers within GPU-aware MPI is not possible on these systems.
-The impact of multiple processes per GPU on the CCL libraries, however, was evaluated on Cray systems.
-Note that NVIDIA MPS was not enabled during benchmarking.
-4.1. MPI Analysis
-All MPI all-reduce tests were analyzed using OpenMPI 5.0.5 and CUDA 12.4.0.
+Thus, each reduction using multiple processes per GPU analyzed on Tuolumne copies to and from another device buffer allocated by the calling process before calling the MPI routines on the local buffer.
+Note that NVIDIA MPS was not enabled during benchmarking on Delta whereas an equivalent implementation by AMD is implicitly enabled on Tuolumne.
+Figure 8: Multi-lane GPU-aware performance on 8 nodes of Delta, varying buffer sizes
+V-A Host Copy Analysis
+All all-reduce tests on Delta were analyzed using OpenMPI 5.0.5 and CUDA 12.4.0 using copies to and from host memory managed by OpenMPI.
 All processes were assigned to the GPU adjacent to their NUMA region.
@@
 All plots contain error bars showing the run to run variation of the algorithms.
-Figure 5 compares the performance of a standard MPI_Allreduce operation against the multi-lane optimization.
+Figure 9: Multi-lane GPU-aware performance of $2^{27}$ floats on Delta, varying node counts
+Figures 8 and 9 compare the performance of a standard MPI_Allreduce operation against the multi-lane optimization.
 These tests are using a single process per GPU, and using all four GPUs available per node on Delta.
@@
 The speedups for the multi-lane algorithm increase with node count, achieving 2.12.1x speedups at $32$ nodes.
-Figure 9. A standard MPI_Allreduce with multiple processes per GPU.
+Figure 10: A standard MPI_Allreduce with multiple processes per GPU on Delta
+Figure 11: The multi-lane all-reduce with multiple processes per GPU on Delta
 Figure 10 shows the performance of a standard MPI_Allreduce when varying the number of processes per GPU from $1$ to $16$.
@@
 Further, these speedups increase with node count through $16$ nodes, yielding over 1.71.7x speedup at $32$ nodes.
-Figure 10. The multi-lane all-reduce with multiple processes per GPU.
-((a)) Tuolumne (RCCL)
-((b)) DeltaAI (NCCL)
-Figure 11. Standard CCL algorithm with a size of $2^{29}$ floats, varying node counts
 Figure 11 displays the performance of the multi-lane algorithm when using multiple processes per GPU.
@@
 Large all-reduce operations across all $4$ GPUs on $8$ nodes yield around a 1.391.39x speedup over the multi-lane algorithm with a single process per GPU.
-Finally, Figure 8 compares the performance of all algorithms, including where the ring algorithm is used in the inter-node stage of the multi-lane all-reduce, comparing $1$ versus $16$ processes per GPU.
+Figure 12: Standard and multi-lane algorithms with multiple processes per GPU on 8 nodes of Delta, varying buffer sizes
+Figure 13: Standard and multi-lane algorithms with multiple processes per GPU of $2^{27}$ floats on Delta, varying node counts
+Finally, Figures 12 and 13 compare the performance of all algorithms, including where the ring algorithm is used in the inter-node stage of the multi-lane all-reduce, comparing $1$ versus $16$ processes per GPU.
 The multi-lane algorithm with $16$ processes per GPU achieves up to 2.42.4x speedup over the standard MPI_Allreduce for large data sizes.
-4.2. CCL Analysis
-This section analyzes the impact of using multiple processes per GPU across CCL all-reduce operations.
-The performance of NCCL is analyzed on a state-of-the-art NVIDIA Grace Hopper architecture, Delta AI, using CUDA 12.3 and NCCL 2.19.3.
-The performance of RCCL is analyzed on one of the fastest AMD architectures, Tuolumne, using ROCm 6.4.0 and RCCL 2.22.3.
+These results agree with our ping-pong microbenchmark on Delta.
+V-B GPUDirect RDMA Analysis
+The performance of GPUDirect RDMA-enabled MPI is analyzed on one of the fastest AMD architectures, Tuolumne, using Cray MPICH 9.0.1 and ROCm 6.4.0.
+For benchmarks using more than one process per GPU, data to be communicated was copied to and from an allocated device buffer before and after MPI communication.
+This is because Cray MPICH does not support buffers shared via ROCm IPC.
+Each process performs an independent all-reduce operation on their assigned subset of the buffer.
+All processes were assigned to the GPU adjacent to their NUMA region.
 Each benchmark was run 3 separate times; error bars in plots show run to run variation.
-Each process performs an independent CCL all-reduce operation on their assigned subset of the buffer.
-As CCL all-reduce operations take a stream, each process per GPU performs this reduction on a separate stream to allow increased parallelism.
-All processes were assigned to the GPU adjacent to their NUMA region.
-Figure 11 shows the performance of a CCL all-reduce at the largest size when varying the number of processes per GPU from 1 to 16 on Tuolumne and DeltaAI.
-On the left featuring Tuolume using RCCL, we see steady speedup as the number of nodes increases with 2 and 4 processes per GPU, yielding speedups of 1.711.71x and 1.591.59x at 8 nodes, respectively.
-These speedups begin to decrease at large scales.
-Using more than 4 processes per GPU does not show any benefit on Tuolumne past 8 nodes with 8 and 16 processes per GPU showing significant slowdown.
-On the right featuring DeltaAI using NCCL, we see substantial speedup at all node counts with 16 processes per GPU, yielding the most significant speedup of 1.771.77x at 16 nodes.
-4 and 8 processes per GPU show speedup only at small scales, while using 2 processes per GPU starts to show benefit at 16 nodes.
-There are many potential reasons for this speedup, including improved network utilization and improved use of NCCL channels.
-To determine if NCCL channels were impacted, the environment variables NCCL_MIN_NCHANNELS and NCCL_MAX_NCHANNELS were set to 4 in order to eliminate the case of low GPU core utilization during memory operations; no differences were observed in timings.
-5. Conclusions and Future Work
+Figure 14: Standard algorithm with multiple processes per GPU on 32 nodes of Tuolumne, varying buffer sizes
+Figure 15: Standard algorithm with multiple processes per GPU of $2^{29}$ floats on Tuolumne, varying node counts
+Figures 14 and 15 show the performance of an all-reduce when varying the number of processes per GPU from 1 to 4 on Tuolumne with each MI300A in SPX mode.
+Figure 14 shows timings on 32 nodes at different sizes of the communication buffer where using multiple processes per GPU shows speedup for the largest buffer sizes.
+Figure 15 shows the timings of the largest buffer size as the number of nodes increases.
+We see steady speedup as the number of nodes increases with 2 and 4 processes per GPU, yielding speedups of 1.171.17x and 1.11.1x at 32 nodes, respectively.
+These speedups begin to decrease at large scales; more significant speedups of up to 1.211.21x are seen at 8 nodes.
+Using more than 4 processes per GPU does not show any benefit on Tuolumne.
+It is noted that a small benchmark was performed on 8 nodes of Tuolumne with the System Direct Memory Access (SDMA) engines of the MI300A both enabled and disabled.
+Consistent with previous findings regarding small-scale inter-APU communication [36], the use of SDMA engines does not appear to impact the advantages provided by using multiple processes per GPU.
+V-B1 MI300A Partitioning Modes
+Figure 16: Standard and multi-lane algorithms with multiple processes per GPU on 32 nodes across the SPX/TPX/CPX modes of Tuolumne, varying buffer sizes
+Figure 17: Standard and multi-lane algorithms with multiple processes per GPU of $2^{28}$ floats across the SPX/TPX/CPX modes of Tuolumne, varying node counts
+The AMD MI300A APUs are also tested in the TPX and CPX partitioning modes.
+Figures 16 and 17 show the speedups of using two processes per GPU for both proposed algorithms against the standard MPI_Allreduce implementation for each partitioning mode  [^1] It is important to note that speedups from each partitioning mode cannot be directly compared since the buffer size per logical GPU is not scaled with the number of logical GPUs..
+Figure 16 shows speedups on 32 nodes of Tuolumne with different buffer size.
+In SPX and CPX modes with either the standand or multi-lane algorithm, respectively, there is a significant advantage to using two processes per GPU for larger buffers.
+Figure 17 shows speedups of both algorithms while the number of nodes increases at the largest buffer size tested across all partitioning modes.
+One of the novel algorithms presented provides speedup in SPX and CPX modes; we see that the multi-lane multi-process algorithm yields speedups of up to 33x over the standard algorithm in CPX mode while the standard multi-process algorithm shows no speedup.
+Performance in TPX mode does not show benefit at scale for either algorithm.
+The standard multi-process algorithm is only advantageous in SPX mode.
+Thus, the multi-lane algorithm benefits from more logical GPUs potentially due to the increased memory locality near each XCD die [3] with smaller communication sizes from the initial call to reduce-scatter.
+The standard algorithm exhibits the opposite, potentially benefiting from less logical GPUs per node, in agreement with the SPX ping-pong microbenchmark on Tuolumne.
+We also observe that our multi-lane algorithm using two processes per GPU in CPX mode scales significantly better than the standard algorithm in any of the partitioning modes; the speedup of our multi-lane algorithm continues to increase as the number of nodes increases.
+Figure 18: The multi-lane all-reduce with multiple processes per GPU on the MI300A partitioning modes on Tuolumne
+The performance of the multi-lane algorithm as the number of nodes increases can be seen in Figure 18.
+We observe that using multiple processes per GPU benefits the multi-lane algorithm in SPX and CPX modes, yielding 1.51.5x and 2.12.1x speedup, respectively, emphasizing that the speedup of the multi-lane algorithm in Figure 17 is due to both the algorithmic structure, the use of multiple processes per GPU, and the fine-grained parallelism offered by CPX partitioning.
+VI Conclusions and Future Work
 In this paper, we proposed to utilize more of the available CPU cores across all GPUs on a node to speed up the GPU-aware all-reduce collective.
-Using an Ampere-based system (Delta) on OpenMPI with our standard and lane-aware approaches, we showed up to an 2.452.45x speedup on large buffers.
-Using RCCL with our standard approach on an El Capitan-based system, we showed a 1.711.71x speedup.
-Using NCCL on a Hopper-based system, we showed a 1.771.77x speedup.
-We conclude that using more of the available CPU cores on modern systems should enhance all-reduce performance at large scales.
+Using an Ampere-based system (Delta) on OpenMPI backed by host copy communication with our standard and lane-aware approaches, we showed up to an 2.452.45x speedup on large buffers.
+Using Cray MPICH backed by GPUDirect RDMA communication with our standard approach on an El Capitan-based system (Tuolumne), we showed a 1.171.17x speedup.
+Across the three partitioning modes of the AMD MI300A on Tuolumne, we showed that the multi-process multi-lane all-reduce is highly performant in CPX mode, yielding a 33x speedup at scale.
 We propose to extensively evaluate the performance characteristics of our multi-process approaches and also extend them to collectives such as the neighborhood all-to-all used in sparse linear systems.
@@
 Emerging systems like Tuolumne feature memory architectures that are shared between the CPU and GPU.
-Future work could potentially examine tradeoffs using CPUs for communication and GPUs for computation, allowing these steps to overlap while minimizing synchronization costs.
+Future work could potentially examine tradeoffs in scientific applications by driving communication with more CPU cores while using GPUs for local computation, allowing these steps to overlap while minimizing synchronization costs.
 Acknowledgments
@@
 This research used the Tuolumne supercomputer at Lawrence Livermore National Laboratory.
-This research used both the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois.
-Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
-Generative AI was used to improve the formatting of data within plots.
+This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois.
+Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
+Generative AI was used to improve formatting of plots.